OpenStudy (anonymous):

WILL FAN AND MEDAL!!!!!! The volume of a cylinder is 24 ft3. The volume of a cube is 54 ft3. What is the difference between the volume of a cone that fits exactly inside the cylinder and the volume of a pyramid that fits exactly inside the cube?

OpenStudy (anonymous):

@mathslover @mathman123 @Hero

OpenStudy (anonymous):

Please help! I'm not sure how to do this

OpenStudy (anonymous):

volume of the cone which will fit into cylinder should have same radius and height so volume of cone = (1/3)*pi*r^2*h = (1/3)*24 = 8

OpenStudy (anonymous):

similarly volume of pyramid will be (1/3)*a*b*c = (1/3)*54 =18

OpenStudy (anonymous):

Thanks. Can you help me with another one?

OpenStudy (anonymous):

yes why not